# A von Neumann algebraic approach to Quantum Theory on curved spacetime

### v3

@@ Abstract.
This manuscript has no associated data.
The authors whose names are listed above declare that they have no conflicts of interest regarding the subject matter or materials discussed in this manuscript.
@@ 1. Introduction
The important point to note here is the fact that $C^{*}$-algebras can be considered as a noncommutative analogue of the family of continuous functions while von Neumann algebras can be considered as a noncommutative counterpart of measurable functions. This fact places von Neumann algebras at the heart of noncommutative integration theory. In our works, [35, 33, 31], we proposed a description of large quantum systems (statistical physics, QFT) based on noncommutative analysis, i.e. on noncommutative integration theory and noncommutative differentiation. Since as noted above (abstract) von Neumann algebras are at the heart of noncommutative integration theory, we shall therefore use abstract von Neumann algebras when engaged in algebraic quantization. We emphasize that the use of abstract von Neumann algebras is not bound in any way to the choice of any particular state. See section 3 for details. In fact as we shall see in section 7 good states which well illustrate emerge from the very geometry of the underlying Lorentzian manifold.
One very obvious advantage of the von Neumann algebraic framework is the uniqueness of the von Neumann algebra tensor product. So the problem of having to decide which tensor product to use never arises. But this raises the question of how the requirement of not demanding a distinguished a priori Hilbert space is to be satisfied. We will demonstrate that the theory of the standard form of von Neumann algebras defines a tensorial category which enables one to incorporate Hilbert spaces into the picture at a local level. As such it fulfils the criterion of not demanding the a priori existence of a distinguished global Hilbert space of states, but nevertheless leaves room for a theory of entanglement at a local level.
In concluding this section we draw attention to some recent papers on this general topic written from a von Neumann algebra perspective. Specifically the following:
@@
    An algebra of observables for de Sitter space by Chandrasekaran, Longo, Penington, Witten [12]. Here the authors describe an algebra of observables for a static patch in de Sitter space, with operators gravitationally dressed to the worldline of an observer. The algebra is a von Neumann algebra of Type $II_{1}$. There is a natural notion of entropy for a state of such an algebra. In their paper the authors also point out the difficulty of isolating a concept of entropy for type III factors. As a contribution to this debate the authors of the present paper wish to point out that in one of their earlier papers [35, Definition 5.3], they identified a class of states who do admit a concept of entropy which proves to be a natural generalisation of von Neumann entropy.
@@ 2. Preliminaries
When $C^{*}$-algebras are in view we shall denote them by $\mathcal{A},\mathcal{B}$, with the symbols $\mathcal{M},\mathcal{N}$ being used to denote von Neumann algebras. Given a faithful normal semifinite weight $\nu$ on a von Neumann algebra, one may of course pass to the GNS-representation $(\pi_{\nu}:\mathcal{M}\to B(H_{\nu}))$ of $\mathcal{M}$} with respect to $\nu$. We will hereafter abbreviate faithful normal semifinite to just fns. With $\eta$ denoting the canonical embedding of the left-ideal $\mathfrak{n}_{\nu}=\{a\in\mathcal{M}\colon\nu(a^{*}a)<\infty\}$ into $H_{\nu}$, one may of course densely define an antilinear closable operator $S_{0}$ on $H_{\nu}$ by the prescription $\eta(a)\to\eta(a^{*})$, $a\in\mathcal{M}$. With $S$ denoting the closure of $S_{0}$, the operator $\Delta=S^{*}S$ is then referred to as the modular operator and the anti-unitary $J$ in the polar decomposition $S=J\Delta^{1/2}$ as modular conjugation. It is well known that the prescription $\sigma_{t}^{\nu}:a\to\pi_{\nu}^{-1}(\Delta^{it}\pi_{\nu}(a)\Delta^{-it})$ $a\in\mathcal{M}$ defines an automorphism group on $\mathcal{M}$} (the so called modular automorphism group) and $j_{\nu}:f\to JfJ$ ($J\in\pi_{\nu}(\mathcal{M})$) an anti-linear *-isomorphism from $\pi_{\nu}(\mathcal{M})$ to $\pi_{\nu}(\mathcal{M})^{\prime}$. On for the sake of ease of notation identifying $\mathcal{M}$ with $\pi_{\nu}(\mathcal{M})$, the natural positive cone $\mathscr{P}_{\nu}$ of $H_{\nu}$ is then defined to be the closure of the set $\{\eta{aj(a)}\colon a\in\mathcal{M}\}$. The content of Haagerup’s theorem regarding the standard form of a von Neumann algebra, is that the quintuple $(\mathcal{M},\pi_{\nu},H_{\nu},j_{\nu},\mathscr{P}_{\nu})$ uniquely identifies the GNS representation of $\mathcal{M}$ up to a spatial *-isomorphism.
@@
The local algebra generated by solutions to the Klein-Gordon equation and indexed by $\mathcal{K}(M,g)$ does turn out to be a quasi-local algebra. Following [3] we may construct this algebra as follows: For each $\mathcal{O}\in\mathcal{K}(M,g)$, one firstly constructs $L^{2}(\mathcal{O})$ where integration is by the volume form on $\mathcal{O}$. The algebra ${\mathcal{A}}(\mathcal{O})$ is then a copy of the $\mathrm{CCR}$ algebra constructed from Weyl operators $W(\psi)$ where the $\psi$ ranges over the solutions of the local Klein-Gordon equation on $\mathcal{O}$, namely $\{\psi\colon\psi=\widetilde{G}_{\mathcal{O}}(f),f\in\mathscr{D}(\mathcal{O},E)\}$. Here we note that the operator $\widetilde{G}_{\mathcal{O}}$ is for every $f\in\mathscr{D}(\mathcal{O},E)\}$ given by $\widetilde{G}_{\mathcal{O}}(f)=G(f_{ext})_{|\mathcal{O}}$ where $f_{ext}$ denotes the element of $\mathscr{D}(M,E)$ obtained by assigning the value 0 to all points $p\in M\backslash\mathcal{O}$. The local algebras of Brunetti, Fredenhagen and Verch [9] are very similar with the main difference being that their Weyl operators are constructed from elements of the set $\{\psi\colon\psi=\widetilde{G}_{\mathcal{O}}(f),f\in C_{0}^{\infty}(\mathcal{O})$. Readers wishing to see an axiomatic description of this approach would be well-advised to consult the papers of Fredenhagen and Rejzner [22], and Buchholz and Fredenhagen [11].
Von Neumann local algebras are typically constructed by passing to a representation by means of a regular state of the $C^{*}$-algebra setting and then taking the double commutant of the representation. At the forefront of such “good” states are the quasi-free states. Representations with these states typically lead to the iconical Araki-Woods factors. See [14] for an account of this process. However in many cases more than just quasi-freeness is needed. Hadamard states have the added advantage that they are related to Quantum Weak Energy Inequalities (QWEIs) which in turn ensure that the existence of “good” Wightman fields at the microscopic level harmonises with observance of the second law of thermodynamics at the macroscopic level. Further physical reasons also bring passive and stationary states into the picture. This topic will be explored in more detail in Part II As noted previously [14] is a good reference for information on quasi-free states and Araki-Woods factors. The paper [30] gives useful information regarding quasi-free, Hadamard and stationary states.
@@ 3. The standard form of von Neumann algebras
We will show that there is a categorical way to incorporate Hilbert spaces into the present framework without going to the extreme of assuming the a priori existence of some global Hilbert space of states. On a philosophical level we agree with the fact that the theory should be based on the principles of AQFT. In the modern approach to canonical quantization this means that to each quantum system there corresponds a $C^{*}$ or von Neumann algebra which encodes the information of that system. For us the basic object is however an abstract von Neumann algebra equipped with some fns weight. With this approach each such abstract von Neumann algebra describing some local region of space-time allows for the construction of a unique Hilbert space associated with that particular algebras by means of the GNS process. This Hilbert space is therefore a product of the system rather than an a priori given object.
@@
For the reader’s convenience, we recall that the equivalent abstract definition of von Neumann algebras identifies these algebras as $C^{*}$-algebra which have a (necessarily unique) predual. In fact such abstract von Neumann algebras are the departure point for the treatment of this topic in both of [45] and [48]. What is crucial for our work is the following well known fact: any von Neumann algebra $\mathcal{M}$ (abstract or concrete) admits a faithful normal weight $\nu$ for which the associated GNS construction yields a faithful standard representation $\pi$ of $\mathcal{M}$, i.e. a normal -isomorphism into the algebra of all bounded linear operators $B(H_{\nu})$ on the GNS Hibert space $H_{\nu}$ with the image $\pi(\mathcal{M})$ being a von Neumann algebra in standard form, see either [49, Theorem IX.1.2 & Definition IX.1.13] or [25, Theorem 10.59].
If one brings the theory of the standard form of a von Neumann algebra into play, one can see that von Neumann algebras in standard form are a category which admit a tensor structure and which does not require an a priori global Hilbert space. Since the von Neumann algebra tensor product is unique there is here no selection of the “correct” tensor product that needs to be made. In addition the tensor structure associated to this category is refined enough to allow for a description of entangled states. We briefly outline these facts.
@@
The standard form of a von Neumann algebra is an abstract characterisation of the uniqueness of the GNS-representation of the pair $({\mathcal{M}},\nu)$ where $\nu$ is a faithful normal semifinite weight on the von Neumann algebra $\mathcal{M}$}. Haagerup proved a very deep theorem essentially showing that any representation of $\mathcal{M}$} which admits objects that mimic the action of $J_{\nu}$ and $\mathscr{P}_{\nu}^{\natural}$ is a faithful copy of the GNS-representation of the pair $({\mathcal{M}},\nu).$ This claim may be made exact with the following definition:
@@
Based on the above theorem it is clear that one may define a category VN-STDFM for which the objects are the quintuples described in Definition 3.1 and the isomorphisms unitary equivalence as described in Theorem 3.2. If both $({\mathcal{M}}_{0},H_{0},J_{0},\mathscr{P}_{0})$ and $({\mathcal{M}}_{1},H_{1},J_{1},\mathscr{P}_{1})$ are in standard form, that means that $({\mathcal{M}}_{0},H_{0},J_{0},\mathscr{P}_{0})$ must be the product of a GNS construction performed with respect to some fns weight $\nu$ on ${\mathcal{M}}_{0}$. A homomorphism from $({\mathcal{M}}_{0},H_{0},J_{0},\mathscr{P}_{0})$ to $({\mathcal{M}}_{1},H_{1},J_{1},\mathscr{P}_{1})$ would then consist of a faithful normal conditional expectation $\mathbb{E}$ satisfying $\nu\circ\mathbb{E}=\nu$ such that $({\mathcal{M}}_{1},H_{1},J_{1},\mathscr{P}_{1})$ is isomorphic to the quadruple produced by the pair $(\mathbb{E}({\mathcal{M}}_{0}),\nu_{|\mathbb{E}({\mathcal{M}}_{0})})$. The structure of these morphisms ensure that any theory constructed using this category will behave well with respect to subtheories.
3.1. VN-STDFM is a tensorial category

### v4

@@ 1. Introduction
The important point to note here is the fact that $C^{*}$-algebras can be considered as a noncommutative analogue of the family of continuous functions while von Neumann algebras can be considered as a noncommutative counterpart of measurable functions. This fact places von Neumann algebras at the heart of noncommutative integration theory. In our works, [38, 36, 34], we proposed a description of large quantum systems (statistical physics, QFT) based on noncommutative analysis, i.e. on noncommutative integration theory and noncommutative differentiation. Since as noted above (abstract) von Neumann algebras are at the heart of noncommutative integration theory, we shall therefore use abstract von Neumann algebras when engaged in algebraic quantization. We emphasize that the use of abstract von Neumann algebras is not bound in any way to the choice of any particular state. See section 3 for details.
One very obvious advantage of the von Neumann algebraic framework is the uniqueness of the von Neumann algebra tensor product. So the problem of having to decide which tensor product to use never arises. But this raises the question of how the requirement of not demanding a distinguished a priori Hilbert space is to be satisfied. We will demonstrate that the theory of the standard form of von Neumann algebras defines a tensorial category which enables one to incorporate Hilbert spaces into the picture at a local level. More specifically this approach will fulfil the criterion of not demanding the a priori existence of a distinguished global Hilbert space of states by relegating Hilbert spaces to a local phenomenon, thereby still leaving room for a theory of entanglement at a local level. It should be noted that the application of the standard form of von Neumann algebras to local algebras is nothing new, although to the best of our knowledge it doesn’t seem to have done in the context of curved spacetime yet. As early as 1984 Doplicher and Longo already demonstrated the utility of these ideas in the context of Minkowski space - see [17].
Our approach in achieving these objectives will be inductive in that we will analyse the behaviour of the local von Neumann algebras constructed using the free Klein-Gordon field with respect to these paradigms and then formalise the emergent structure into an axiomatic framework. In concluding this section we draw attention to some recent papers on this general topic written from a von Neumann algebra perspective. Specifically the following:
@@
    An algebra of observables for de Sitter space by Chandrasekaran, Longo, Penington, Witten [12]. Here the authors describe an algebra of observables for a static patch in de Sitter space, with operators gravitationally dressed to the worldline of an observer. The algebra is a von Neumann algebra of Type $II_{1}$. There is a natural notion of entropy for a state of such an algebra. In their paper the authors also point out the difficulty of isolating a concept of entropy for type III factors. As a contribution to this debate the authors of the present paper wish to point out that in one of their earlier papers [38, Definition 5.3], they identified a class of states who do admit a concept of entropy which proves to be a natural generalisation of von Neumann entropy.
@@ 2. Preliminaries
When $C^{*}$-algebras are in view we shall denote them by $\mathcal{A},\mathcal{B}$, with the symbols $\mathcal{M},\mathcal{N}$ being used to denote von Neumann algebras. Given a faithful normal semifinite weight $\nu$ on a von Neumann algebra, one may of course pass to the GNS-representation $(\pi_{\nu}:\mathcal{M}\to B(H_{\nu}))$ of $\mathcal{M}$} with respect to $\nu$. We will hereafter abbreviate faithful normal semifinite to just fns. With $\eta$ denoting the canonical embedding of the left-ideal $\mathfrak{n}_{\nu}=\{a\in\mathcal{M}\colon\nu(a^{*}a)<\infty\}$ into $H_{\nu}$, one may of course densely define an antilinear closable operator $S_{0}$ on $H_{\nu}$ by the prescription $\eta(a)\to\eta(a^{*})$, $a\in\mathfrak{n}_{\nu}\cap\mathfrak{n}^{*}_{\nu}$. With $S$ denoting the closure of $S_{0}$, the operator $\Delta=S^{*}S$ is then referred to as the modular operator and the anti-unitary $J$ in the polar decomposition $S=J\Delta^{1/2}$ as modular conjugation. It is well known that the prescription $\sigma_{t}^{\nu}:a\to\pi_{\nu}^{-1}(\Delta^{it}\pi_{\nu}(a)\Delta^{-it})$, where $a\in\mathcal{M}$, defines an automorphism group on $\mathcal{M}$} (the so called modular automorphism group) and $j_{\nu}:f\to JfJ$ ($J\in\pi_{\nu}(\mathcal{M})$) an anti-linear *-isomorphism from $\pi_{\nu}(\mathcal{M})$ to $\pi_{\nu}(\mathcal{M})^{\prime}$. For the sake of ease of notation we will identify $\mathcal{M}$ with $\pi_{\nu}(\mathcal{M})$. The natural positive cone $\mathscr{P}_{\nu}$ of $H_{\nu}$ may then be defined to be the closure of the set $\{\pi_{\nu}(a)J_{\nu}(\eta(a))\colon a\in\mathfrak{n}_{\nu}\cap\mathfrak{n}^{*}_{\nu}\}$ ([52, p. 146]). In the case where $\nu$ is a state, this formula simplifies to $\{\eta(aj(a))\colon a\in\mathcal{M}\}$. The content of Haagerup’s theorem regarding the standard form of a von Neumann algebra, is that the quintuple $(\mathcal{M},\pi_{\nu},H_{\nu},j_{\nu},\mathscr{P}_{\nu})$ uniquely identifies the GNS representation of $\mathcal{M}$ up to a spatial *-isomorphism.
@@
The local algebra generated by solutions to the Klein-Gordon equation and indexed by $\mathcal{K}(M,g)$ does turn out to be a quasi-local algebra. Following [3] we may construct this algebra as follows: For each $\mathcal{O}\in\mathcal{K}(M,g)$, one firstly constructs $L^{2}(\mathcal{O})$ where integration is by the volume form on $\mathcal{O}$. The algebra ${\mathcal{A}}(\mathcal{O})$ is then a copy of the $\mathrm{CCR}$ algebra constructed from Weyl operators $W(\psi)$ where the $\psi$ ranges over the solutions of the local Klein-Gordon equation on $\mathcal{O}$, namely $\{\psi\colon\psi=\widetilde{G}_{\mathcal{O}}(f),f\in\mathscr{D}(\mathcal{O},E)\}$. Here we note that the operator $\widetilde{G}_{\mathcal{O}}$ is for every $f\in\mathscr{D}(\mathcal{O},E)\}$ given by $\widetilde{G}_{\mathcal{O}}(f)=G(f_{ext}){\upharpoonright}{\mathcal{O}}$ where $f_{ext}$ denotes the element of $\mathscr{D}(M,E)$ obtained by assigning the value 0 to all points $p\in M\backslash\mathcal{O}$. The local algebras of Brunetti, Fredenhagen and Verch [9] are very similar with the main difference being that their Weyl operators are constructed from elements of the set $\{\psi\colon\psi=\widetilde{G}_{\mathcal{O}}(f),f\in C_{0}^{\infty}(\mathcal{O})$. Readers wishing to see an axiomatic description of this approach would be well-advised to consult the papers of Fredenhagen and Rejzner [23], and Buchholz and Fredenhagen [11].
Von Neumann local algebras are typically constructed by passing to a representation by means of a regular state of the $C^{*}$-algebra setting and then taking the double commutant of the representation. At the forefront of such “good” states are the quasi-free states. Representations with these states typically lead to the iconical Araki-Woods factors. See [14] for an account of this process. However in many cases more than just quasi-freeness is needed. Hadamard states have the added advantage that they are related to Quantum Weak Energy Inequalities (QWEIs) which in turn ensure that the existence of “good” Wightman fields at the microscopic level harmonises with observance of the second law of thermodynamics at the macroscopic level. Further physical reasons also bring passive and stationary states into the picture. This topic will be explored in more detail in Part II. As noted previously [14] is a good reference for information on quasi-free states and Araki-Woods factors. The paper [32] gives useful information regarding quasi-free, Hadamard and stationary states.
@@ 3. The standard form of von Neumann algebras
We will show that there is a categorical way to incorporate Hilbert spaces into the present framework without going to the extreme of assuming the a priori existence of some global Hilbert space of states. On a philosophical level we agree with the fact that the theory should be based on the principles of AQFT. In the modern approach to canonical quantization this means that to each quantum system there corresponds a $C^{*}$ or von Neumann algebra which encodes the information of that system. For us the basic object is however an abstract von Neumann algebra equipped with some fns weight. With this approach each such abstract von Neumann algebra describing observables associated with some local region of space-time allows for the construction of a unique Hilbert space associated with that particular algebra by means of the GNS process. This Hilbert space is therefore a product of the system rather than an a priori given object.
@@
For the reader’s convenience, we recall that the equivalent abstract definition of von Neumann algebras identifies these algebras as $C^{*}$-algebra which have a (necessarily unique) predual. In fact such abstract von Neumann algebras are the departure point for the treatment of this topic in both of [48] and [51]. What is crucial for our work is the following well known fact: any von Neumann algebra $\mathcal{M}$ (abstract or concrete) admits a faithful normal weight $\nu$ for which the associated GNS construction yields a faithful standard representation $\pi$ of $\mathcal{M}$, i.e. a normal -isomorphism into the algebra of all bounded linear operators $B(H_{\nu})$ on the GNS Hibert space $H_{\nu}$ with the image $\pi(\mathcal{M})$ being a von Neumann algebra in standard form, see either [52, Theorem IX.1.2 & Definition IX.1.13] or [27, Theorem 10.59].
At this point there are two related categorical options for taking the theory further. We know from Maharam’s theorem that the category of abelian von Neumann algebras equipped with some faithful normal semifinite trace is equivalent to the category of localizable measure algebras (see [27, Chapter 1]). In a quantum setting the pairings $(\mathcal{M},\nu)$ where $\nu$ is a faithful normal semifinite weight may therefore be regarded as noncommutative analogues of localizable measure algebras, and which we may therefore refer to as von Neumann measure algebras (cf. Definition 3.5). It is this very fact that underlies the rationale behind using von Neumann algebras as a formalism for Quantum information Theory. Here two such pairings, $(\mathcal{M}_{0},\nu_{0})$ and $(\mathcal{M}_{1},\nu_{1})$, are equivalent if there is a *-isomorphism $\iota$ from $\mathcal{M}_{0}$ onto $\mathcal{M}_{1}$ such that $\nu_{0}=\nu_{1}\circ\iota$. A measure algebraic approach would therefore be based on a categorical analysis of these pairings. However, as we shall shortly see, as far as quantum information theory is concerned an approach which more strongly emphasises the GNS construction may be more beneficial as those are the tools which facilitate a description of entanglement. The two approaches are interlinked in that equivalent pairings $(\mathcal{M},\nu)$ will produce standard forms that are copies of each other, but they are not synonymous. When bringing the theory of the standard form of a von Neumann algebra into play, it quickly becomes clear von Neumann algebras in standard form are also a category which admit a tensor structure and which does not require an a priori global Hilbert space, since the form itself encodes the “canonical” Hilbert space for that von Neumann algebra. On a related note, we observe that since the von Neumann algebra tensor product is unique there is, in contrast to the $C^{*}$ context, here no selection of the “correct” tensor product to be made. In addition the tensor structure of standard forms of von Neumann algebras is refined enough to allow for a description of entangled states, and therefore by extension a refined theory of quantum information theory. We briefly outline these facts.
@@
The standard form of a von Neumann algebra is an abstract characterisation of the uniqueness of the GNS-representation of the pair $({\mathcal{M}},\nu)$ where $\nu$ is a faithful normal semifinite weight on the von Neumann algebra $\mathcal{M}$}. Just as in classical measure where replacing a given measure with an equivalent measure will leave most of the measure theoretic constructs intact, different faithful normal semifinite weights may produce a GNS space which in all aspects is a copy of the first. Haagerup proved a very deep theorem giving expression to this fact by essentially showing that any representation of $\mathcal{M}$} which admits objects that mimic the action of $J_{\nu}$ and $\mathscr{P}_{\nu}^{\natural}$ is a faithful copy of the GNS-representation of the pair $({\mathcal{M}},\nu).$ This claim may be made exact with the following definition:
@@
Definition 3.3.
On the basis of the above theorem we define a category VN-SF for which the objects are the quintuples described in Definition 3.1. The isomorphisms for this category are the spatial *-isomorphisms described in Theorem 3.2.
Next let $(\mathcal{N},\nu_{\mathcal{N}})$ and $(\mathcal{M},\nu_{\mathcal{M}})$ be pairings for which $\mathcal{N}$ is for some projection $e\in{\mathcal{M}}$ a sub-von Neumann algebra of $e{\mathcal{M}}e$ and $\nu_{{\mathcal{N}}}=\nu_{\mathcal{M}}{\upharpoonright}{{\mathcal{N}}}$ to $\mathcal{N}$} is still semifinite. Let $e_{({\mathcal{M}}:{\mathcal{N}})}$ be the projection of $H_{\nu_{{\mathcal{M}}}}$ (the GNS space of the pairing $(\mathcal{M},\nu_{\mathcal{M}})$) onto the subspace $H_{{\mathcal{N}}}$ generated by $\{\eta_{\nu}(a):a\in\mathfrak{n}_{\nu_{{\mathcal{N}}}}\cap\mathfrak{n}^{*}_{\nu_{{\mathcal{N}}}}\}$. It can easily be seen that $H_{{\mathcal{N}}}$ is a copy of $H_{\nu_{{\mathcal{N}}}}$. We then say that the standard form of a pairing $(\mathcal{N},\nu_{\mathcal{N}})$ embeds into the standard form of the pairing $(\mathcal{M},\nu_{\mathcal{M}})$ if it is of the form $({\mathcal{N}},\pi_{\nu_{{\mathcal{M}}}}{\upharpoonright}{{\mathcal{N}}},H_{\mathcal{N}},J_{\nu_{{\mathcal{M}}}}{\upharpoonright}{H_{{\mathcal{N}}}},\mathscr{P}_{\nu_{{\mathcal{N}}}})$ where $\mathscr{P}_{\nu_{{\mathcal{N}}}}\subset\mathscr{P}_{\nu_{{\mathcal{M}}}}$.
Morphisms in the category VN-SF are the simply defined to be any combination of embeddings and isomorphisms.
@@
It is clear from [27, §10.5] that subalgebras $\mathcal{N}$ of some von Neumann algebra $\mathcal{M}$ equipped with a faithful normal semifinite weight $\nu$ for which there either exists a normal conditional expectation $\mathbb{E}:\mathcal{M}\to\mathcal{N}$ with $\nu\circ\mathbb{E}=\nu$, or a projection $e$ in the centralizer $\mathcal{M}_{\nu}$ for which $\mathcal{N}=e\mathcal{M}e$ with $\nu{\upharpoonright}{e\mathcal{M}e}$ still semifinite, provide particular examples of the standard form embeddings described above.
The category that best suits our needs is in fact the subcategory of the above for which the objects are standard forms of pairings $({\mathcal{M}},\nu)$ of von Neumann algebras and an associated fns weight (referred to as von Neumann measure algebras in standard form), since this subcategory retains the benefits of access to standard forms, but more strongly emphasises the integration theory aspect.
Definition 3.5.
The category VN-MA-SF is defined as follows:
Objects:
The objects consist of quintuples $({\mathcal{M}},\nu,H_{\nu},J_{\nu},\mathscr{P}_{\nu})$ where $\mathcal{M}$} is a von Neumann algebra equipped with a faithful normal semifinite weight, and $(H_{\nu},J_{\nu},\mathscr{P}_{\nu})$ respectively the Hilbert space, modular conjugation and natural positive cone generated by performing the GNS construction for the pair $({\mathcal{M}},\nu)$.
Morphisms:
Two objects $({\mathcal{M}}_{1},\nu_{1},H_{\nu_{1}},J_{\nu_{1}},\mathscr{P}_{\nu_{1}})$ and $({\mathcal{M}}_{2},\nu_{2},H_{\nu_{2}},J_{\nu_{2}},\mathscr{P}_{\nu_{2}})$ are said to be isomorphic if there exists a *-isomorphism $\alpha$ from ${\mathcal{M}}_{1}$ onto ${\mathcal{M}}_{2}$ such that $\nu_{2}\circ\alpha=\nu_{1}$. (Recall that we have already noted that if such a *-isomorphism exists, it will induce a unitary operator from $H_{\nu_{1}}$ to $H_{\nu_{2}}$ which identifies the triples $(H_{\nu_{1}},J_{\nu_{1}},\mathscr{P}_{\nu_{1}})$ and $(H_{\nu_{2}},J_{\nu_{2}},\mathscr{P}_{\nu_{2}})$.) We say that $({\mathcal{M}}_{1},\nu_{1},H_{\nu_{1}},J_{\nu_{1}},\mathscr{P}_{\nu_{1}})$ embeds into $({\mathcal{M}}_{2},\nu_{2},H_{\nu_{2}},J_{\nu_{2}},\mathscr{P}_{\nu_{2}})$ if for some projection $e\in{\mathcal{M}}_{2}$, ${\mathcal{M}}_{1}$ is a subalgebra of $e{\mathcal{M}}_{2}e$, $\nu_{1}=\nu_{2}{\upharpoonright}{\mathcal{M}}_{1}$, $H_{\nu_{1}}\subseteq H_{\nu_{2}}$, $J_{\nu_{1}}=J_{\nu_{2}}{\upharpoonright}H_{\nu_{1}}$ and $\mathscr{P}_{\nu_{1}}\subseteq\mathscr{P}_{\nu_{2}}$. Morphisms are then defined to be combinations of isomorphisms and embeddings.
The following Lemma provides criteria under which the condition $J_{\nu_{1}}=J_{\nu_{2}}{\upharpoonright}H_{\nu_{1}}$ pertains.
Lemma 3.6.
Let $\mathcal{M}$} be a von Neumann algebra equipped with an fns weight $\nu$ and let $\mathcal{N}$} be a von Neumann subalgebra for which $\nu{\upharpoonright}{\mathcal{N}}$ is still semifinite. Then the GNS Hilbert space of the pair $({\mathcal{N}},\nu{\upharpoonright}{\mathcal{N}})$ corresponds to the subspace $H_{{\mathcal{N}}}=\overline{\{\eta_{{\mathcal{M}}}(a):a\in\mathfrak{n}_{\nu}\cap\mathfrak{n}^{*}_{\nu}\cap{\mathcal{N}}\}}$ of $H_{\nu}=H_{{\mathcal{M}}}$. If $H_{{\mathcal{N}}}$ is an invariant subspace of the modular conjugation operator $J_{{\mathcal{M}}}$ for $({\mathcal{M}},\nu)$, then under this identification $J_{{\mathcal{M}}}{\upharpoonright}H_{{\mathcal{N}}}$ is exactly the modular conjugation operator $J_{{\mathcal{N}}}$ for the pair $({\mathcal{N}},\nu{\upharpoonright}{\mathcal{N}})$, in which case $\mathscr{P}_{{\mathcal{N}}}\subseteq\mathscr{P}_{{\mathcal{M}}}$
Proof.
It is clear that the closable densely defined anti-linear operator
$S_{0}^{{\mathcal{N}}}:\eta_{{\mathcal{M}}}(a)\mapsto\eta_{{\mathcal{M}}}(a^{*})\mbox{ for all }a\in\mathfrak{n}_{\nu}\cap\mathfrak{n}^{*}_{\nu}\cap{\mathcal{N}}$
corresponding to the pair $({\mathcal{N}},\nu{\upharpoonright}{\mathcal{N}})$, is a restriction of the closable densely defined anti-linear operator
$S_{0}^{{\mathcal{M}}}:\eta_{{\mathcal{M}}}(a)\mapsto\eta_{{\mathcal{M}}}(a^{*})\mbox{ for all }a\in\mathfrak{n}_{\nu}\cap\mathfrak{n}^{*}_{\nu}.$
If momentarily we regard the closed extension $S^{{\mathcal{N}}}$ of $S_{0}^{{\mathcal{N}}}$ as a map into $H_{{\mathcal{M}}}$, it is then clear from [52, Lemma VI.1.4] that $S^{{\mathcal{M}}}$ is an extension of $S^{{\mathcal{N}}}$. That is $S^{{\mathcal{M}}}{\upharpoonright}\mathrm{dom}(S^{{\mathcal{N}}})=S^{{\mathcal{N}}}$. Since $H_{{\mathcal{N}}}$ is an invariant subspace of $J_{{\mathcal{M}}}$, the operator $J_{{\mathcal{M}}}$ restricts to an anti-unitary on $H_{{\mathcal{N}}}$. Let us write $\widehat{J}_{{\mathcal{M}}}$ for this restriction. We will then have that $\widehat{J}_{{\mathcal{M}}}S^{{\mathcal{N}}}$ is a closed operator on $H_{{\mathcal{N}}}$ for which we have that
$\widehat{J}_{{\mathcal{M}}}S^{{\mathcal{N}}}=J_{{\mathcal{M}}}S^{{\mathcal{M}}}{\upharpoonright}\mathrm{dom}(S^{{\mathcal{N}}})=\Delta^{1/2}_{{\mathcal{M}}}{\upharpoonright}\mathrm{dom}(S^{{\mathcal{N}}}).$
In view of the fact that $\mathrm{dom}(S^{{\mathcal{N}}})\subset\mathrm{dom}(S^{{\mathcal{M}}})$, it is palpably clear that $\langle\xi,\Delta^{1/2}_{{\mathcal{M}}}\xi\rangle\geq 0$ for all $\xi\in\mathrm{dom}(S^{{\mathcal{N}}})$ and hence that $\Delta^{1/2}_{{\mathcal{M}}}{\upharpoonright}\mathrm{dom}(S^{{\mathcal{N}}})$ is a positive operator. By construction $S^{{\mathcal{N}}}=\widehat{J}_{{\mathcal{M}}}(\Delta^{1/2}_{{\mathcal{M}}}{\upharpoonright}\mathrm{dom}(S^{{\mathcal{N}}}))$ is then a polar decomposition of $S^{{\mathcal{N}}}$. Since $\Delta^{1/2}_{{\mathcal{N}}}$ has dense range, the uniqueness of the polar decomposition now ensures that $J_{{\mathcal{N}}}=\widehat{J}_{{\mathcal{M}}}$ as was required. (Although we shall not need this, we also have that $\Delta^{1/2}_{{\mathcal{N}}}=\Delta^{1/2}_{{\mathcal{M}}}{\upharpoonright}\mathrm{dom}(S^{{\mathcal{N}}})$.)
The final claim is now an obvious consequence of the definition of the natural positive cone [27, Definition 8.40]. ∎
3.1. VN-SF and VN-MA-SF are tensorial categories